Run Planning Meeting Notes, April 10 - April 16, 2025

From GlueXWiki
Jump to: navigation, search

↑ Back to Run Coordination Meetings:Spring2025#Runplan


Wednesday April 16

  1. Last 24 hours
    1. Trigger commissioning
    2. Ran Compton configuration for ECAL voltages
  2. Shift summaries
    1. Day Swing Owl
  3. Reports
    1. trig_client restarted
    2. Helicty board in readout and working Please test this data.
    3. Run starts now in hdrun log
    4. Beam excursion history
    5. Atmospheric pressure sensor noisy when solenoid ramping
    6. New FCAL voltages loaded
    7. Compton events after ECAL HV adjustment
    8. DIRC scalers restored
  4. Issues
    1. Solenoid ramp down Solenoid ramp down ramping instructions
    2. Tagger voltage issues
    3. Beam
      1. Continue to have excursions
  5. Plans
    1. RF recovery, beam projected by 9pm
    2. Hall A switching to 5th pass and 50 uA in about 10 days
    3. Finish aligning diamonds. Will align 1 thin and 1 thick at Hovanes convenience.
  6. Access today
    1. Hall D
      1. Provakar (test area)
      2. Lubomir +2 (TRD)
    2. Tagger ~10am
      1. Voltage crate power cycle
      2. Hovanes
      3. INC

Tuesday April 15

  1. Last 24 hours
    1. Ran Compton configuration for ECAL voltages
    2. Some trigger development work
    3. Started CVC HV Scan
  2. Shift summaries
    1. Day Swing Owl
  3. Reports
    1. TAGM bias voltages updated configs updated Microscope scalers are much smoother.
    2. CDC/FDC HV scan analysis Scan done last week.
    3. New ECAL voltages loaded
  4. Issues
    1. Beam
      1. Still get regular excursions on the AC
      2. Beam was missing the collimator on Day Shift for a period due to tuning for Hall A. Hall b also saw issues.
      3. Starts in arcAPathlength drift to blame "Currently the locks are keeping the beam on target ain hall D, but if it starts drifting extensively we should consider checking/adjusting pathlength both in ArcA and earlier."
      4. Small beam energy oscillation 1 MeV over 15 minutes. Likely a sick cavity.
    2. Helicity Board status
    3. DIRC scalers not counting?
  5. Plans
    1. Make FCAL voltage change today
    2. RF recover Wednesday Day Shift
    3. Hall A will go to 5th pass in 2 weeks time. They are starting early and ending early to get the highest priority data done first. That means our high polarization period will come earlier than expected.

Monday April 14

  1. Last 3 days:
    1. Friday
      1. Access during Friday Day Shift. PLC stuffChill water hose
      2. Loaded Tagger production voltages from 2023 which turns off all taggers below the Microscope.
      3. Raised beam current to 200 nA for production.
    2. Saturday
      1. Diamond Alignment JD80-212 position XY scan position XY scan position chosen
      2. Started TOF HV scan. Procedure
    3. Sunday
      1. TOF HV Scan completed.
      2. Production
  2. Shift summaries
    1. Friday: Day Swing Night Club
    2. Saturday: Owl Day Night Club Swing
    3. Sunday: Owl Day Shift Swing
    4. Monday Owl
  3. Reports
    1. TOF recabling successful
  4. Beam
    1. Slow ramp turned off
    2. No lock to AC Hall D orbit
    3. Tief did some investigations.
    4. Position variability
    5. Position excursions after beam trip.
    6. Lock indicator on CSS general screen.
    7. Mode beam stuff
  5. Issues
    1. ECAL IOC
    2. TAGH Voltage signal mixup
    3. ESAD documentation issue
    4. TOF Voltage takes long but can be worked around
    5. halldweb down
    6. random triggers missing
  6. Plans
    1. CDC HV scan
    2. RF recovery Wednesday

Friday April 11

Ran BCAL+ECAL+PS trigger during Thursday Day shift. Early Swing, trigger was modified to remove BCAL prescale factor. Rate went up to 22 kHz.

Accelerator Schedule:

  • Day Shift: RF recovery 8-11 am, and pass change HB->5 until Swing.
  • If RF recovery goes long then pass change moves to Saturday.
  • Next maintenance day, Wednesday 16 April

Reports

Shift summaries DaySwingNight ClubOwl

Issues

Thursday April 10

Ran BCAL+ECAL+PS trigger overnight. A lot of data taken but significant DAQ live time issues. The triggers are not timed in together.

Reports

Shift summaries Day shiftOwl shiftSwing shift

Issues